OpenStudy (highschoolmom2010):

For a regular n-gon: a. What is the sum of the measures of its angles? b. What is the measure of each angle? c. What is the sum of the measures of its exterior angles, one at each vertex? d. What is the measure of each exterior angle? e. Find the sum of your answers to parts b and d. Explain why this sum makes sense.

OpenStudy (highschoolmom2010):

a. (n-2)180

OpenStudy (primeralph):

B. divide your answer in (a) by n

OpenStudy (primeralph):

C. 360

OpenStudy (primeralph):

D. 360/n

OpenStudy (primeralph):

do E by yourself.

OpenStudy (highschoolmom2010):

no value was given for n though

OpenStudy (bahrom7893):

n depends on the number of angles.

OpenStudy (highschoolmom2010):

no value wsa given for n so how do i do e.

OpenStudy (bahrom7893):

(n-2)*180/n + 360/n = ( (n-2)*180 + 360 )/n =(180n-360+360)/n = 180n/n = 180

OpenStudy (highschoolmom2010):

so n in the end just equals 180?

OpenStudy (bahrom7893):

no, no matter what your n is, you end up with 180.